OpenStudy (lovely_lovely):

State whether each extreme in the data set is an outlier. A. The lower extreme is an outlier, but the upper extreme is not. B. The upper extreme is an outlier, but the lower extreme is not. C. Both extremes are outliers. D. Neither extreme is an outlier.

OpenStudy (anonymous):

You have to find interquartile range, which is Q3 - Q1 Q3 = 41 Q1 = 36 IQR = Q3 - Q1 = 41 - 36 = 5 The outliers will be any values below 36 - 1.5 * 5 = 36 - 7.5 = 28.5 or above 41 + 1.5 * 5 = 41 + 7.5 = 48.5 Therefore, your outliers are : 25 and 45 (C)

OpenStudy (anonymous):

can i get a gold medal?

OpenStudy (lovely_lovely):

the answer is c

OpenStudy (mathmale):

SwagGirl: This is the second time within a few minutes that I've seen you asking for a medal. Would you kindly stop. Read the Code of Conduct to see what the rules are in this case.

OpenStudy (anonymous):

Sorry @mathmale